OpenStudy (anonymous):

a^2-10a+25=0 Please show step by step. Thank You.

OpenStudy (anonymous):

a^2 -5a -5a+25=0 a(a-5)-5(a-5)=0 (a-5)(a-5)=0 or (a-5)^2=0 or a=5,5

OpenStudy (anonymous):

Thank you but MyMathLab program is saying that 5,5 is not correct

OpenStudy (anonymous):

(a - 5)^2 = 0 a - 5 = 0 a = 5

OpenStudy (anonymous):

because (-5)^2 = 25 and 2(-5) = -10

OpenStudy (anonymous):

ah ... okay. Thank you!

OpenStudy (anonymous):

uw!